# Reference Coverage Analysis of OpenAlex compared to Web of Science and Scopus

### v2

@@ 3.1 Total and Source Reference Coverage
Table 3, in a naïve averaging of the source reference count, leaves OpenAlex looking comparatively poor at 7.6 references per record to the 16.9 or 18.7 of WoS and Scopus (and well behind the other databases’ average total reference count). However, when restricting to the 2015-2022 corpus shared by the three databases, OpenAlex proves competitive with a higher average source reference count than both WoS and Scopus. The fact that results vary greatly depending on the underlying corpus definition could be explained by the fact that OpenAlex comprises of many–possibly less scientific–publications with comparatively short reference lists. When focusing on the comparison of the average total reference counts between WoS and Scopus, it initially appears that Scopus outperforms WoS, however when considering records marked as articles they perform more comparably. This trend continues when observing the Shared Corpus and the Shared Corpus with references from 1996 to 2022. Notably here the difference between the source reference count and total reference count decreases as the restrictions are added. The results suggest that Scopus still has a small disadvantage due to its initial indexing start in 1996. Consequently, the slight advantage for OpenAlex is reversed when references are restricted to reference publication years 1996-2022, with Scopus outperforming OpenAlex, and WoS performing worst – however, differences are very small.
@@
Table 3: Comparison of the reference coverage available in each database, including the reference counts from the database providers, and our computed counts
@@ 3.2 Discrepancies between Reference Counts and Reference Data
The detected discrepancies raise issues with both proprietary and open source bibliometric databases and should be considered when working with the references, and likely merits a deeper analysis.
@@
3.3 Metadata Coverage
@@
Figure 2 highlights the metadata coverage analysis results between OpenAlex and the two proprietary databases, WoS and Scopus, within the Shared Corpus. The x-axis represents OpenAlex, while the y-axis corresponds to WoS (left) and Scopus (right). The points represent the percentage coverage of the relevant indicator per journal.
@@
Figure 2: Plots of the coverage of metadata between OpenAlex and Web of Science and Scopus
@@
The results indicate that OpenAlex depicts a different pattern compared to WoS and Scopus in terms of abstracts (Figure 2a), with the two proprietary databases having a higher overall availability of abstracts. In total, over 92% of the articles in WoS and Scopus have abstract information, compared to a 87% coverage of abstracts in OpenAlex. In contrast, the ORCID coverage is more comprehensive in OpenAlex (Figure 2b). The proportion of articles in OpenAlex with at least one ORCID present is 92%, and the proportion of articles with at least one ORCID in WoS is 16% and in Scopus 32%. However, upon inspection we discovered that OpenAlex performs a generous disambiguation of authors, resulting in a high ORCID coverage, which may explain this large difference. In particular, some authors with Chinese names were linked to more than 10,000 publications. The distribution of open access information is more linear (Figure 2c), with a tendency slightly in favour of OpenAlex, suggesting an indexing lag of Unpaywall’s open access status information in the WoS and Scopus data. The proportion of open access information in all three databases is around 49%.
3.4 Discussion
@@
This study demonstrates the source reference coverage in OpenAlex to be comparable to that in WoS and Scopus for comparatively newer records which lie in the intersection of all three databases, both in general and when restricting to references from 1996 onwards. On the one hand, this could be seen as an indicator of good quality bibliometric core data. On the other hand, OpenAlex does not have the highest internal coverage, although it is by far the largest database, so it would actually be plausible that higher proportions of the referenced publications are themselves part of the database. In this respect, the Scopus coverage policy seems to be a bit more effective. However, one possible factor could also be that a comparatively poorer reference-matching algorithm misses a noticeable amount of actual source references. From table 3 it can be inferred that within the Shared Corpus, there are on average 6.4 to 6.2 references captured in the total reference count by WoS and Scopus (respectively) that OpenAlex does not capture in its source reference count. The fact that OpenAlex does not systematically include non-source references, as well as complete reference strings, limits the flexibility of using and exploring the data source: It does not allow researchers or bibliometric centers to apply their own reference matching algorithms or to analyse non-source references as such.
@@
The study also revealed data errors in Scopus and OpenAlex. The reported figures for reference counts in Scopus do not correspond to the actual numbers of references in the database, and OpenAlex is inconsistent in its handling of references as it does not systematically comprise all non-source references, but references to deleted source items. Similarly, we note that all databases, to a different degree, comprise cases where DOIs refer to multiple records–cf. [5]. We believe it merits further study and caution when replicating these computations.
@@
Our analysis reveals that OpenAlex demonstrates a particularly high level of coverage for ORCID in comparison to WoS and Scopus. Over 90% of articles in OpenAlex had been assigned at least one ORCID. However, we have observed that this percentage is somewhat excessive. Upon inspection, we discovered that in some case ORCIDs were assigned to more than 10,000 records in our corpus, suggesting issues with OpenAlex’s author disambiguation method. Notably, Chinese author names were among those most affected.
@@ 4 Limitations and Outlook
In a more extensive setting, an in-depth comparison of source and non-source references of each publication in a sample between the databases could provide indications of the extent to which the detected smaller differences can be explained by different coverage profiles or strengths and weaknesses of the matching algorithms. A possible extension of our main methodological setting could analyse the internal coverage with respect to the disciplinary level and address the question to what extent OpenAlex has a better (or worse) coverage of non-English, regionally-oriented journals which might be relevant to some Arts & Humanities and Social Sciences subjects, for example, and do not easily fulfill WoS curation criteria.
When studying ORCID availability, it must be noted that we did not check for the availability for all co-authors, but just if there was at least one ORCID present per article. It is important to conduct further analysis to confirm whether the author names and ORCIDs are accurately matched, given the observed phenomenon of a single ORCID being erroneously attributed to tens of thousands of articles. If this is not the case then this may demonstrate the ongoing challenge of author name disambiguation, in particular in the case of Chinese names.
As discussed in Section 2.2, some DOIs were found to have duplicate records assigned to them in each of the three databases, requiring us to deselect the 39,481 records from 2015-2022 which lay in the intersection of the three databases and had more than one record associated with the DOI in one of the databases from our Shared Corpus. A more detailed examination of duplicate DOIs may be merited, in particular with respect to Scopus (as demonstrated by Table 2.) Similarly investigations into the distribution of duplicated DOI by record type between each database may be recommended for future research.
@@
Funding This work was funded by the Federal Ministry of Education and Research via funding numbers: 16WIK2301B / 16WIK2301E, The OpenBib project. We acknowledge support by Federal Ministry of Education and Research, Germany under grant number 01PQ17001, the Competence Network for Bibliometrics.
@@
Conflicts of interests/Competing interests
Philipp Mayr, the co-author of this paper, has a conflict of interest because he serves on the editorial board of the journal Scientometrics.

### v3

@@ 3.1 Total and Source Reference Coverage
Table 3, in a naïve averaging of the source reference count, leaves OpenAlex looking comparatively poor at 7.6 references per record to the 16.9 or 18.7 of WoS and Scopus (and well behind the other databases’ average total reference count). However, when restricting to the 2015-2022 corpus shared by the three databases, OpenAlex proves competitive with a higher average source reference count than both WoS and Scopus. The fact that results vary greatly depending on the underlying corpus definition suggests that OpenAlex comprises of many publications with comparatively short reference lists which are not contained by the WoS or Scopus. When focusing on the comparison of the average total reference counts between WoS and Scopus, it initially appears that Scopus outperforms WoS, however when considering records marked as articles they perform more comparably. This trend continues when observing the Shared Corpus and the Shared Corpus with references from 1996 to 2022. Notably here the difference between the source reference count and total reference count decreases as the restrictions are added. The results suggest that Scopus still has a small disadvantage due to its initial indexing start in 1996. Consequently, the slight advantage for OpenAlex is reversed when references are restricted to reference publication years 1996-2022, with Scopus outperforming OpenAlex, and WoS performing worst – however, differences are very small.
@@
Table 3: Comparison of the reference coverage available in each database, including the reported reference counts from the database providers, the pre-calculated source reference counts from FIZ-Karlsruhe, and our computed counts
@@ 3.2 Discrepancies between Reference Counts and Reference Data
The detected discrepancies between both proprietary and open source bibliometric databases should be considered when working with OpenAlex for bibliometric analyses - as averages of reference counts may differ significantly if the databases are not judiciously curated. We believe this discrepancy likely merits a deeper analysis in OpenAlex as new versions are released.
@@
3.3 Metadata by Journal
Continuing the analysis of OpenAlex, WoS and Scopus, we then broke down the data by journal. Firstly in Figure 2, we compared the counts of source references in each journal in the WoS and Scopus to those in OpenAlex, spotting a fairly similar distribution in the two comparisons. Comparing these to Figure 3, we observe that the greater density under the $y=x$ line indicates OpenAlex is on average identifying slightly more source references in some journals, but the lesser density above the line indicates OpenAlex significantly undercounts on some journals as compared to WoS and Scopus.
Figure 2: Scatter diagrams of the count of source references per journal between OpenAlex and the Web of Science and Scopus.
Figure 3: Scatter diagrams of the count of source references per journal between OpenAlex and the Web of Science and Scopus.
@@ 3.3 Metadata by Journal
Figure 4 then highlights the metadata coverage analysis results between OpenAlex and the two proprietary databases, WoS and Scopus, within the Shared Corpus. The x-axis represents OpenAlex, while the y-axis corresponds to WoS (left) and Scopus (right). The points represent the percentage coverage of the relevant indicator per journal.
@@
Figure 4: Scatter diagrams of the coverage of metadata per journal between OpenAlex and the Web of Science and Scopus.
The results indicate that OpenAlex depicts a different pattern compared to WoS and Scopus in terms of abstracts (Figure 4a), with the two proprietary databases having a higher overall availability of abstracts. In particular we note that there are concentrations near the top of the plot, indicating that the proprietary databases have full access to abstract information where OpenAlex has either partial or no access to this information.
Observing the top right of Figure 4a, we see that the majority of journals reside in this area; in total, over 92% of the articles in WoS and Scopus have abstract information, compared to a 87% coverage of abstracts in OpenAlex, which implies a density in the top right hand corner which is not well indicated in the scatter diagrams. Furthermore we see below the $y=x$ line that for some journals OpenAlex has a higher abstract coverage.
@@
In contrast, the ORCID coverage is more comprehensive in OpenAlex (Figure 4b). The proportion of articles in OpenAlex with at least one ORCID present is 92%, and the proportion of articles with at least one ORCID in WoS is 16% and in Scopus 32%. However, upon inspection we discovered that OpenAlex performs a generous disambiguation of authors, resulting in a high ORCID coverage. In particular, some authors with Chinese names were observed to be linked to more than 10,000 publications.
The distribution of open access information is more similar between the databases (Figure 4c), with a tendency slightly in favour of OpenAlex, suggesting an indexing lag of Unpaywall’s open access status information in the WoS and Scopus data. The proportion of open access information in all three databases is around 49%.
Viewing Figure 4d shows that the availability of funding information on articles is better represented in WoS and Scopus than in OpenAlex. Notably, funding information associated with articles in over 4,100 journals can only be found in WoS and Scopus, which could indicate a lack of provision of funding information by some scholarly publishers for open databases such as OpenAlex and Crossref.
4 Discussion
@@
This report demonstrates the source reference coverage in OpenAlex to be comparable to that in WoS and Scopus for comparatively newer records which lie in the intersection of all three databases, both in general and when restricting to references from 1996 onwards. On the one hand, this can be seen as an indicator of good quality bibliometric core data. On the other hand, OpenAlex does not have the highest internal coverage, although it is by far the largest database, so it would actually be plausible that higher proportions of the referenced publications are themselves part of the database. In this respect, the Scopus coverage policy seems to be a bit more effective. However, one possible factor could also be that a comparatively poorer reference-matching algorithm misses a noticeable amount of actual source references.
The vastly greater corpus of document records in OpenAlex, compared to WoS and Scopus, raises the question of what this additional content is, which is covered by OpenAlex but by neither established commercial provider. Our findings demonstrate what this content is not: it is not that part of the scientific literature which is referenced by items within WoS or Scopus. If that were the case, we would have found that OpenAlex internal reference coverage clearly exceeding that of the other to data sources in the Shared Corpus, because more references cited by those publications would be indexed by OpenAlex, but not WoS and Scopus. The substantially larger differences between the mean source reference counts of OpenAlex and the other two databases, if the entire databases and not the fixed comparison corpus are taken, also shows that the publications that are only in OpenAlex and not in the other two databases pull down the mean values due to their lower reference counts. They must therefore represent a different publication spectrum or have a significantly lower data quality. In any case, this suggests that OpenAlex should be limited to a core corpus if comparability of bibliometric analyses based on OpenAlex to WoS and Scopus is desired.
From Table 3 it can be inferred that within the Shared Corpus, there are on average 6.4 to 6.2 references captured in the total reference count by WoS and Scopus (respectively) that OpenAlex does not capture in its source reference count. The fact that OpenAlex does not yet systematically include non-source references, as well as complete reference strings, limits the flexibility of using and exploring the data source: It does not allow researchers or bibliometric centers to apply their own reference matching algorithms or to analyse non-source references as such.
@@
The study also revealed data errors in Scopus and OpenAlex. The reported figures for reference counts in Scopus do not correspond to the actual numbers of references in the database, and OpenAlex is inconsistent in its handling of references as it does not systematically comprise all non-source references, but references to some deleted source items. Similarly, we note that all databases, to a different degree, comprise cases where DOIs refer to multiple records–cf. [8]. We believe it merits further study and caution when replicating these computations. Another study by [11], in preprint at time of writing, indicates that data errors have been discovered in the ”is_retracted” field of OpenAlex for publications between 22 December 2023 and 19 March 2024, further highlighting the volatility of the metadata quality in OpenAlex.
In summary, from an internal reference coverage perspective, OpenAlex as a source for citation data for studies of contemporary scientific output, is on par with commercial databases when restricted to a core corpus of publications similar to that of WoS and Scopus. However, its utility is hampered by not yet providing full cited reference data.
@@
Our analysis reveals that OpenAlex demonstrates a particularly high level of coverage for ORCID in comparison to WoS and Scopus. Over 90% of articles in OpenAlex had been assigned at least one ORCID. However, we have observed that this percentage is somewhat excessive. Upon inspection, we discovered that in some cases ORCIDs were assigned to more than 10,000 records in our corpus, suggesting issues with OpenAlex’s author disambiguation method. The authorships and author records have also been subject to updates by OpenAlex since data collection, including cleaning of author strings, syncing 17.9 million work records with Crossref, removing 3.9 million empty author records (authors with no works assigned to them) and updating author information fields.
In conclusion, our analysis of the metadata by journal highlights data collection and curation challenges for OpenAlex, having to collate information from both bibliometric and non-bibliometric sources [^9] https://help.openalex.org/hc/en-us/articles/24397285563671-About-the-data requires OpenAlex to perform disambiguation and standardisation between data sources, both challenging tasks, as well as deal with legal constraints in collecting and publishing academic works – for example, the copyright of the abstracts. These challenges likely differ from those of WoS and Scopus in their collection and curation, but the similarity of the figures plotting OpenAlex against WoS and OpenAlex against Scopus demonstrate a stark difference between OpenAlex, and WoS and Scopus. Therefore we currently recommend caution when utilising OpenAlex for scientometric studies due to the volatility and data quality issues discussed earlier in this section.
@@ 5 Limitations and Outlook
We restate that our data is representative of late 2023, with the hitherto noted volatility of OpenAlex in the time since, this report may not be representative of the state of OpenAlex, and also the Web of Science and Scopus, at time of publishing.
@@
In a more extensive setting, an in-depth comparison of source and non-source references of each publication in a sample between the databases could provide indications of the extent to which the detected smaller differences can be explained by different coverage profiles or strengths and weaknesses of the matching algorithms. A possible extension of our main methodological setting could analyse the internal coverage with respect to the disciplinary level and address the question to what extent OpenAlex has a better (or worse) coverage of non-English, regionally-oriented journals which might be relevant to some arts & humanities and social sciences subjects, for example, and do not easily fulfil WoS curation criteria.
When studying ORCID availability, it must be noted that we did not check for the availability for all co-authors, but just if there was at least one ORCID present per article. It is important to conduct further analysis to confirm whether the author names and ORCIDs are accurately matched, given the observed phenomenon of a single ORCID being erroneously attributed to tens of thousands of articles. If this is not the case then this may demonstrate the ongoing challenge of author name disambiguation in bibliographic databases.
As discussed in Section 2.1, some DOIs were found to have duplicate records assigned to them in each of the three databases, requiring us to deselect the 39,481 records from 2015-2022 which lay in the intersection of the three databases and had more than one record associated with the DOI in one of the databases from our Shared Corpus. A more detailed examination of duplicate DOIs may be merited, in particular with respect to Scopus (as demonstrated by Table 2 and agreeing with findings in reported in [27] on Scopus.) Similarly investigations into the distribution of duplicated, missing or incorrect DOI by record type between each database may be recommended for future research.
Therefore as highlighted in [27] and [18], where procedures for merging and deduplicating bibliographic datasets are explored, we note that DOIs are not perfect identifiers for combining datasets and often, due to duplicated, missing or incorrect DOIs, lead to the requirement for more demanding title and author name comparisons for generation of better quality datasets for future study.
Since the data collection effort of this study, at least 151 million references have been added to OpenAlex, as of the May 30th 2024 snapshot of OpenAlex reportedly expanding the number of references by 7.61% compared to the April 25th 2024 snapshot. This and other ongoing efforts to improve data quality and expand data availability in OpenAlex indicate the need for future similar studies, and highlight the volatility of the database.
@@
Funding This work was funded by the Federal Ministry of Education and Research via funding numbers: 16WIK2301B / 16WIK2301E, The OpenBib project [23]. We acknowledge support by Federal Ministry of Education and Research, Germany under grant number 01PQ17001, the Competence Network for Bibliometrics.
@@
Prior Public Release
The findings presented in this paper were initially reported in a preprint on ArXiv [3]. That version includes a longer and more extensive version of the results.